OpenStudy (anonymous):

PLEASE HELP!!!! A ball is thrown vertically from the ground at an initial velocity of 100 ft per sec. How long does it take for the ball to come back to the ground? (Hint: Use the formula h = -16t^2 + v0t + h0, where h is the initial height in feet, t is the time is seconds, v0 is the initial velocity in feet per second, and h0 is the initial height in feet.) Please also explain how you'd get the answer to this problem. Thanks! =)

jimthompson5910 (jim_thompson5910):

h = -16t^2 + v0t + h0 h = -16t^2 + 100t + 0 ... Note: v0 = 100 and h0 = 0 h = -16t^2 + 100t So the height equation is h = -16t^2 + 100t The ball will come back to the ground when h = 0, so plug it in and solve for t h = -16t^2 + 100t 0 = -16t^2 + 100t -16t^2 + 100t = 0 -4t(4t - 25) = 0 -4t = 0 or 4t - 25 = 0 t = 0/(-4) or 4t = 25 t = 0 or t = 25/4 Ignore the first solution as that's the point in time where the ball takes off So the ball will come back down when t = 25/4 or t = 6.25 seconds

OpenStudy (anonymous):

Wow! Now, this is a thorough explanation, if ever I saw one! :) I actually do understand it, now! I was so confused. Thank you so much - this cleared up everything!

jimthompson5910 (jim_thompson5910):

that's great, glad I could be of help

OpenStudy (anonymous):

Maybe you should go into teaching/tutoring math - you're really good at it!

jimthompson5910 (jim_thompson5910):

lol I already am a tutor, but thank you for the compliment

OpenStudy (anonymous):

Oh, well that makes sense, then. Well, you're the best tutor I've ever had, so far. :) Keep up the great work!

jimthompson5910 (jim_thompson5910):

That means a lot to me, thank you.

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

and to add a bonus, you're smiling in the face of math (while others are probably frowning)...that's good lol

OpenStudy (anonymous):

Ha, ha - I like how you phrased that! Makes me feel like I'm standing on the battlefield of math, as a hero. :)

jimthompson5910 (jim_thompson5910):

lol that's one good way to look at it

OpenStudy (anonymous):

:)